OpenStudy (anonymous):

Express the concentration of a 0.0390 M aqueous solution of fluoride, F–, in mass percentage and in parts per million. Assume the density of the solution is 1.00 g/mL.

OpenStudy (australopithecus):

so, So density doesn't matter ppm = mg/L or Kg solution Mass % = Mass of Solute/Mass of solution again just come up with an amount of solution im going to use 1L 0.0390 M = x moles/1L x = 0.0390Moles of F- 0.0390mol*18.998403g/mol = 0.7409g PPM: 1mg = 1000g 0.7409g = 740.9mg 740.9mg/1L = 740.9mg Mass Percent: density 1g/mL thus 1g = 1mL thus 1L = 1kg thus 1L = 1000g so, (0.7409g/1000g)*100 = 0.07409%_(m/m)